Find three consecutive integers such that the sum of the second and third exceeds 1/2 of the first by 33.
Answer by CubeyThePenguin(3113)

consecutive integers: (x-1), x, (x+1)

x + (x+1) = 1/2 * (x-1) + 33
2x + 1 = 1/2 * (x-1) + 33
4x + 2 = (x-1) + 66
4x + 2 = x + 65
3x = 63
x = 21

The numbers are 20, 21, and 22.